OpenStudy (anonymous):

When the price of a dozen eggs is raised by $2.4, 15 fewer eggs can be bought with $60. (a) Find the original price of an egg. (b) How many eggs can be bought after raising the price?

OpenStudy (anonymous):

@satellite73 would you kindly help me?

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

lets call the price per egg, which we don't know, as \(x\) then if the price goes up by $2.4 a dozen, then it goes up by \(2.4\div 12=0.2\) per egg

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i am sure there is an easier way to do this but i will continue

OpenStudy (anonymous):

if we call the number of eggs bought for $60 \(y\) then before the price increase we have \[xy=60\] and after the price increase we have \[(x+.2)(y-15)=60\]

OpenStudy (anonymous):

y is ?

OpenStudy (anonymous):

something tells me there is an easier way, but i can't see it at the moment

OpenStudy (anonymous):

\(y\) is the number of eggs you can buy for \(\$60\) before the price increase

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that is why \(xy=60\)

OpenStudy (anonymous):

or we can say \(x=\frac{60}{y}\) that is the original price per egg

OpenStudy (anonymous):

that turns the equation \[(x+.2)(y-15)=60\] into \[(\frac{60}{y}+0.2)(y-15)=60\] and you can solve this for \(y\)

OpenStudy (anonymous):

so the equations are: xy=60 (x+0.2)(y-15)=60 ?

OpenStudy (anonymous):

or for that matter you can solve \[(x+0.2)(\frac{60}{x}-15)\]

OpenStudy (anonymous):

wait a second, let me calculate

OpenStudy (anonymous):

i have an answer, i am not sure i like it, but i have one

OpenStudy (anonymous):

lol my calculation is wrong, i got 40 something

OpenStudy (anonymous):

i got \(x=4\) and \(y=75\)

OpenStudy (anonymous):

meaning originally you paid $4 per egg, which seems rather steep then again so does a $2.40 price per dozen increase are you sure it was not $.24 ?

OpenStudy (anonymous):

the answer of part a is $0.8

OpenStudy (anonymous):

it is $2.40 increase, or \(\$.24\) per dozen

OpenStudy (anonymous):

can you show me your working steps? maybe i got my steps wrong

OpenStudy (anonymous):

you mean in solving the equation, or how i got the equation?

OpenStudy (anonymous):

solving the equation

OpenStudy (anonymous):

ooohhhh!!!! i made a mistake you are right, it is \(.8\)

OpenStudy (anonymous):

the answer is 0.8, but i can't get it

OpenStudy (anonymous):

solve \[(x+0.2)(\frac{60}{x}-15)=60\]

OpenStudy (anonymous):

i got 43.33333.....

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

okay we can work out the steps

OpenStudy (anonymous):

 is it?